OpenStreetMap Foundation, Operations Meeting* - Agenda & draft minutes
Wednesday 16 June 2021 18:00 London time
Location: Video room at https://osmvideo.cloud68.co
* Please note that this was not strictly an Operations Working Group (OWG) meeting.
- Tom Hughes (OWG)
- Grant Slater (OWG)
- Paul Norman (OWG)
- Guillaume Rischard (OSMF board)
- Michal Migurski
Minutes by Dorothea Kazazi.
Related to 2020-09-09 Grant [Topic: AWS] Talk to OpenAerial Map/HOT.
Paul talked to HOT about OpenAerial Map. We can give them geotiff and can produce tiles from it.
Switches: ordered 8 10G fiber models, Grant to receive them. GBP 39 /each.
Should get some SFP+ DAC cables (copper ones, offer lowest latency and higher reliability) - need passive ones.
Action item: Grant to get cables for switches.
Patch panel: ordered.
Travel: Grant could probably go there towards the end of July (barring any Covid-19 related restrictions by UK).
- Payment: Delay and partial payment (just VAT) - won't get them by 1st July.
- Shipment: to Cambridge. Payment was partial, so address can be changed.
15' Michal had to leave.
HP pizza boxes from German reseller
Order: ~ 7 servers.
Action item: Paul to ask for discount regarding order of 7 HP pizza boxes from German reseller.
Press announcement for RSS/torrent functionality
Action item: Paul to review the draft text.
Monthly Operations Working Group reports
- Paul goes through the OPS and Chef repositories to see changes.
- Data Working Group does them every 3 months and it is some work.
Suggestion: Send big news to the IRC channel and collect them afterwards.
Any other business
2021-06-02 Grant and Guillaume to finalise the model of switches. Switches to go to the Dublin data center. was: Grant to price up 10Gb Switches options for review and decision. [[Topic: 10G Switches]# 2021-06-16 Done. 2021-06-02 Guillaume to email Equinix Dublin about their recent requirement (related to patch panel) and find out about any additional costs (e.g. crossconnect). [Topic: Dublin Move]# 2021-06-16 Done. 2021-06-02 OWG to look at Paul's proposed specs for ordering DB and planet servers for Dublin in next 24 hrs. [Topic: Dublin Move]# 2021-06-16 Done.
- 2021-06-02 Paul to work out numbers for QGIS sponsoring. [Topic: QGIS] # 2021-06-16 To be done as part of Paul's SotM2021 talk in July, so next update after the talk has been recorded.
- 2021-06-02 Paul to make clear to QGIS that if they don't built a rate limit in the software, they'll get auto-blocked. Bulk geocoding is strongly discouraged. [Topic: QGIS] # 2021-06-16 Assigned to Paul.
- 2021-05-19 Grant to give Twitter credentials to Paul
(was: 2021-05-05 Grant to check/fix GroupTweet for osm_tech Twitter account)# 2021-06-02 pending. 2021-05-19 Grant to price up 10Gb Switches options for review and decision. [Topic: 10GB Switches]# 2021-06-16 Done.
- 2021-05-05 Grant to email Toby from WikiMedia Foundation and suggest chating to MapTiler. [Topic: Wikimedia] # 2021-06-02 & 2021-06-16 pending.
- 2021-03-24 Paul to create ticket related to API PostgreSQL update [Topic: API PostgreSQL update] # 2021-06-02 pending.
- 2021-03-24 Tom to report back on TimescaleDB again at next meeting. [Topic: Reportage] [was: 2021-01-13 Tom to evaluate TimescaleDB] [Topic: Longer term metric retention] # 2021-04-21 SSD Disk Failing in US # 2021-05-19 decision to leave on the agenda. # 2021-06-02 & 2021-06-16 nothing new.
- 2021-02-24 Grant to install a Discourse instance to get us started. [Topic: Discourse] # 2021-04-21 on the agenda. # 2021-05-19 & 2021-06-02 pending.
- 2021-01-13 OWG to send message to the servers we want to keep. [Reportage. Existing CDN servers] # 2021-03-24 Three servers stopped talking to us (shenron, naga and one more) # 2021-05-19, 2021-06-02 & 2021-06-16 pending.
- 2021-01-13 Grant to wipe thorns and the 3 other machines [AMS] [Topic: Longer term metric retention] # 2021-05-19 pending # 2021-06-02 Ramoth data drives wiped - decision: Grant to do final wipe of Ramoth and leave it until next site visit. Discussion about 16G DDR3.
- 2021-01-13 Paul to create ticket with Equinix to scrap the wiped thorns and the other 3 machines [Topic: Longer term metric retention]. # 2021-06-16 waiting for wipe (above)
- 2020-12-02 Grant to develop some thoughts on what is next for us using AWS. [Topic: AWS] # 2021-05-19 & 2021-06-02 pending, 2021-06-16 postponed for a few weeks.
- 2020-11-04 OWG to work out tile log archival and deletion policy at later stage. [Topic: Commercial CDN] # 2021-03-24 & 2021-05-19 deferred to future point.
- 2020-10-21 Paul to write to Discourse ticket and email the board [Topic: Discourse].
- 2020-09-09 Grant [Topic: AWS] Speak to AWS person about going ahead with open data program with official OSM S3 bucket. # 2021-05-19 & 2021-06-02 pending.
- 2020-09-09 Grant [Topic: AWS] Talk to OpenAerial Map/HOT. # 2021-05-19 & 2021-06-02 pending. # 2021-06-16 Paul talked to HOT about OpenAerial Map. We can give them geotiff and can produce tiles from it.
- 2020-08-12 Michal to try to rekindle excitement about people helping with imagery (on dev channel/imagery channel or Slack). # 2020-08-26 No progress.
- 2020-07-29 Grant to enable background sync to AWS S3. [Topic: Ironbelly] # 2020-08-12, 2020-08-26 & 2021-06-02 Manually run, automated scripting to be added. # 2021-05-19 Grant to run the script again.
- 2020-07-29 Grant to check with Wiki Admins on hCaptcha (reCaptcha replacement). [Topic: Wiki reCaptcha issue] https://github.com/openstreetmap/operations/issues/454 # 2020-08-12 hCaptcha people reached out and happy to help. Blocker on Mediawiki 1.35 being released in August. # 2021-05-19 blocker removed. # 2021-06-02 pending.
- 2020-07-15 Paul and|]] Grant to quote up a server to replace errol/kessie. [Topic: Replacement of Errol/Kessie]. # 2020-08-12 A new person in OWG asked to do Errol. Need to replace it at some point - at UCL. # 2021-05-19 & 2021-06-02 pending.
- 2020-07-01 Paul to create a ticket about solutions to reduce incoming comms. [Topic:Revision of acceptable use policy to reduce incoming comms] # 2021-05-19 decision to leave the action item open. # 2021-06-02 discussion about priority for account deletion.
- 2020-07-01 Grant to work out some of the questions for an online form as a solution to reduce incoming comms. [Topic: Revision of acceptable use policy to reduce incoming comms] 2020-08-12 need to think about the reply # 2021-05-19 decision to leave the action item open.
- 2020-06-04 Paul to update the Github ticket "Adding API key support for tile.osm.org" https://github.com/openstreetmap/operations/issues/342
- 2020-04-10 Grant to work out a table of different data bits, work out how they are backed up and what can be potentially improved. [Topic: High Availability / Redundancy of OpenStreetMap.org (and primary services)] # 2021-05-19 decision to leave the action item open. # 2021-06-02 pending.
Wednesday 30 June 2021 18:00 London time
Operations meetings are currently being held every 2 Wednesdays, at 18:00 London time.
Online calendar showing the OPS meetings.